2y(x^2 - y + x) dx + (x^2 - 2y) dy = 0
1
Expert's answer
2021-10-18T04:30:47-0400

"M(x, y)=2y(x^2-y+x), M_y=2x^2-4y+2x"

"N(x, y)=x^2-2y, N_x=2x"


"\dfrac{M_y-N_x}{N}=\dfrac{2x^2-4y+2x-2x}{x^2-2y}=2"

"\mu(x)=e^{\int 2dx}=e^{2x}"

"2e^{2x}y(x^2 - y + x) dx +e^{2x} (x^2 - 2y) dy = 0"

"M_y=2e^{2x}(x^2-2y+x)"

"N_x=e^{2x}(2(x^2-2y)+2x)"

"M_y=2e^{2x}(x^2-2y+x)=N_y"

"u(x, y)=\int e^{2x} (x^2 - 2y)dy+g(x)"

"=e^{2x}x^2y-e^{2x}y^2+g(x)"

"\dfrac{\partial u}{\partial x}=2e^{2x}x^2y+2e^{2x}xy-2e^{2x}y^2+g'(x)"

"=2e^{2x}y(x^2 - y + x)"

"g'(x)=0=>g(x)=-C"

"e^{2x}x^2y-e^{2x}y^2-C=0"

"e^{2x}x^2y-e^{2x}y^2=C"